# GlucoLens: Explainable Postprandial Blood Glucose Prediction from Diet and Physical Activity

### v1

@@ IV-C Alternative Feature Search Methods
Another dimension to explore could be to find possible feature sets other than the five options presented in Fig. 3. One way could be to perform an exhaustive search within the features. The problem with that approach would be the computational cost. As we have 31 features, we would have $2^{31}$ possibilities which would be infeasible to explore. However, we have done an additional experiment to find effective feature subsets. Among all the models, our random forest (RF) was the best model so far with an NRMSE of 0.123. We limited the number of features by restricting the number of leaf nodes in RF models to 24, 48, and 96. The 48-leaf-node limit outperformed the others, achieving an NRMSE of 0.121, surpassing our previous best result of 0.123. This approach offers a more rational feature selection method, as limiting leaf nodes helps identify features with higher information gains, reduces overfitting, and improves generalization.
V Conclusion
@@
In this study, we have analyzed different diet and physical activity features to develop an interpretable ML system for estimating postprandial blood glucose parameters. We believe that our tools will be helpful to people diagnosed with or at risk of diabetes to better regulate their condition and avoid unwanted outcomes. Based on the experiments with different variations of models and feature sets, our GlucoLens system with Random Forest-based backbone outperforms the baseline models by a margin of 12 to 21%. Additionally, with the help of correlation analysis, we established the relationship between PAUC and the risk of hyperglycemia, as well as the features leading to an increase or decrease in the risk of hyperglycemia.
Our proposed GlucoLens system achieves a 74% accuracy on hyperglycemia prediction. Moreover, it recommends different interventions to avoid hyperglycemia based using diverse counterfactual explanations. Based on the explanations, we showed that a regulated diet and higher physical activity can avoid hyperglycemia.

### v2

@@ 4.3 Alternative Feature Search Methods
We also explored another avenue of research to find possible feature sets other than the five options presented in Table 1. One way could be to perform an exhaustive search within the features. The problem with that approach would be the computational cost. As we have 31 features, we would have $2^{31}$ possibilities which would be infeasible to explore. However, we have done an additional experiment to find effective feature subsets. Among all the models, our random forest (RF) was the best model so far with an NRMSE of 0.123. For finding alternative subsets, we limited the number of features by restricting the number of leaf nodes in RF models to 24, 48, and 96. The 48-leaf-node limit outperformed the others, achieving an NRMSE of 0.121, surpassing our previous best result of 0.123. This approach offers a more rational feature selection method, as limiting leaf nodes helps identify features with higher information gains, reduces overfitting, and improves generalization.
5 Conclusion
@@
In this study, we have developed an interpretable ML system for estimating postprandial blood glucose parameters by analyzing different diet and physical activity features. We believe that our tools will be helpful to people diagnosed with or at risk of diabetes to better regulate their condition and avoid unwanted outcomes. We developed a multi-sensor-based system, GlucoLens, that is based on advanced data processing for creating multimodal feature sets and machine learning regressors and classifiers, and large language models for hyperglycemia prediction. Based on the experiments with different variations of models and feature sets, our GlucoLens AUC prediction system in its best configuration outperforms the comparison models by a margin of 10% to 27%. Moreover, it achieves a 73.3% accuracy and an F1 score of 0.716 on hyperglycemia prediction. Finally, it recommends different interventions to avoid hyperglycemia based using diverse counterfactual explanations. Based on the explanations, we showed that a regulated diet and higher physical activity can avoid hyperglycemia.
[…]